Solve the equation for x. Do not include "x =" in your answer.
3+squareroot sign over 2x+4 = 7

Found 2 solutions by solver91311, edjones:
Answer by solver91311(12126)





Add -3 to both sides



Square both sides



You should be able to handle it from here.


John


Answer by edjones(7311)

3+sqrt(2x+4)=7
sqrt(2x+4)=4
2x+4=16
2x=12
x=6
.
Ed